ABX Corp. started their business this year and purchased the following inventory: 15,000 unites on January @ $10 each, 25,000 units in February @ $12 each, 35,000 units in March @ 12 each, 50,000 units in April @ $14 each, 20,000 units in May @ $15 each, 30,000 units in June @ $ 16 each, and 25,000 units in July @ $18 each. During the year they sold 140,000 units of their product. Calculate the value of the ending inventory utilizing the Average Cost, LIFO and FIFO methods.
Answer #1

Ending Inventory

= Units purchased- Units sold

= 200,000 - 140,000

= 60,000

Average cost: $ 840,000

January 15000 10 150000
February 25000 12 300000
March 35000 12 420000
April 50000 14 700000
May 20000 15 300000
June 30000 16 480000
July 25000 18 450000
Total 200,000 2,800,000
Average cost Ending inventory

= 2,800,000 / 200,00

= 14

= 60,000 * 14

= 840,000

LIFO: $ 1,005,000

July 25000 18 450000
June 30000 16 480000
May 5000 15 75000
Ending inventory 60,000 1,005,000

FIFO: $ 690,000

January 15000 10 150000
February 25000 12 300000
March 20000 12 240000
Ending inventory 60,000 690,000